I tried this problem, but just can't solve it!
To avoid hitting any rocks below, a cliff diver jumps up and out. The equation h= -16t(the t is squared) + 4t + 26 decribes her height, h, in feet t seconds after jumping. Find the time at which she returns to a height of 26 feet.
I tried:
26 = -16t(squared) + 4t =26, but can't figure out how to solve

Answer by stanbon(48551)

Subtract 26 from both sides to get:
-16t^2+4t=0
-4t(4t-1)=0
t=0 or t=1/4
At time 0 and at time 1/4 sec. she is at the 26 ft level.
Cheers,
Stan H.